# Boosting Adversarial Transferability for Hyperspectral Image Classification Using 3D Structure-invariant Transformation and Intermediate Feature Distance

### v1

@@ III-A Preliminary
In this study, we refer to the known model used to train and generate adversarial examples as the substitute model, denoted as $f$. The input HSIs are denoted as $x\in X\subset{R^{C\times H\times W}}$, where $H$ and $W$ represent the height and width, respectively, and $C$ is the number of spectral bands. The ground-truth label of $x$ is $y$. The operation of transforming data is denoted as $T\left(\cdot\right)$, and the adversarial example is $x_{adv}\in X\subset R^{C\times H\times W}$. The goal of this study is to find a small perturbation $\delta\subset R^{C\times H\times W}$ that satisfies the constraint condition and enables the adversarial example $x_{adv}=x+\delta$ to maximally mislead the model’s predictions. The loss function is denoted as $L\left(\cdot\right)$ (typically using the cross-entropy loss), so the process of generating adversarial examples is naturally achieved by maximizing the loss function, and it can be expressed as:
${x_{adv}}=\arg\max L\left(f\left(T\left(x_{adv}\right)\right),y\right),s.t.{\left\|\delta\right\|_{p}}\leq\varepsilon$ (1)
Typically, the $l_{p}$-norm is used to constrain the perturbation $\delta$ (in our work, the $l_{1}$-norm is utilized).
@@ III-B Overall Process
The pipeline of our method is illustrated in Fig. 1, The transferability of adversarial examples is enhanced through two key steps: random spatial-spectral block transformation and feature distancing loss. Before feeding the perturbed adversarial examples into the substitute model, $n$ copies are created. These $n$ copies are randomly divided along both the spatial and spectral dimensions. Subsequently, each block will randomly apply a data augmentation method to generate a set of augmented transformed images, which will serve as input for the alternative model. To guide perturbation updates, we integrate intermediate feature distance information between the original and adversarial examples with output layer classification loss to formulate the total loss. Specifically, we first enlarge the original features of the original examples and compute the variance of each channel. The variances of the channels are used to calculate channel weights, which are then incorporated into the weighted Chi-squared distance between the original and adversarial features. The auxiliary classification loss is combined with the feature distancing loss as the total loss function. Because the transformed images may differ significantly from the original ones, the calculation of gradients could introduce instability. To address this, we compute the average gradient across multiple transformed images. In this way, during the iterative optimization process, the perturbation will be updated in the direction of maximizing the feature distance and the prediction distance between the original and adversarial examples, thereby enhancing transferability. The details are introduced below. Algorithm 1 outlines the perturbation generation process integrated with the baseline method MI-FGSM [25].
III-C Block Random Transformation
Data augmentation methods can mitigate the issue of adversarial examples overfitting to the substitute model. Therefore, to enhance the transferability of HSIs, input transformation is essential. Unlike some previous methods, we also consider the structural information of the image. Specifically, given an image $x$, we randomly divide it into $a\times a$ blocks, and the relative relationships among these anchors are the structure of the image, where each anchor is the center of a image block. This structure captures important semantic information for human recognition. Under the premise of preserving the image structure, the transformation of image blocks does not impair recognition but instead generates more diverse images, which is beneficial for improving transferability. On hyperspectral datasets, if we do not disrupt the overall relative structure of the image and only perform block-wise transformations in the spatial dimension, we find that the adversarial examples generated in this way improve the attack transfer success rate compared to baseline methods. However, when compared to the spatial attention maps across various models, the spectral attention maps exhibit more significant differences[33]. Different classification models focus on distinct discriminative spectral bands, so considering spectral neighborhood information may improve transferability.
@@ III-C Block Random Transformation
Figure 2: (a) Schematic of image division. The division parameters in both the spatial and spectral dimensions are set to 2, resulting in 2×2×3 blocks; (b) Schematic of image transformation. Taking spatial transformation as an example, the image is divided into 2×2 blocks, with different transformations applied to each block as shown.
Therefore, in order to further improve the diversity of HSIs, we make full use of the redundancy of the spectral bands in HSIs. First, a set of images is obtained by copying the adversarial examples. On the basis of following the idea of structural invariance, each image is randomly partitioned into $a\times a$ blocks in the spatial dimension and $b$ blocks in the spectral dimension, resulting in $a\times a\times b$ blocks (the division diagram is shown in Fig. 2. Then transform it block by block, randomly select different transformation methods to apply to the image block, and finally get a set of transformed images (the transformation diagram is shown in Fig. 2).
@@
  Input: HSI example $x$, surrogate model $f$, loss function $L$, perturbation budget $\delta$, step size $\alpha$, decay factor $\mu$, momentum term $m_{i}$, gradients $g_{i}$, number of copies $N$, iteration times $I$.
@@
3:     Construct a set $X^{adv}_{aug}$ of $N$ transformed images using $T()$:
4:            $X^{adv}_{aug}=T(x+\delta_{i})^{*}N$
@@
6:     Calculate the feature distancing loss according to (4);
7:     Calculate the auxiliary loss according to (5);
@@
10:     Calculate the average gradient of $N$ copies:
11:           $g_{i}^{ave}=\frac{1}{N}\left(\sum_{j=0}^{N}(\nabla_{\delta}L(x_{j}^{adv},x))\right),(x_{j}^{adv}\in X^{adv}_{aug})$
@@
16:     Update adversarial examples: $x_{i+1}^{adv}=\text{clip}_{x}\{x+\delta_{i}\}$
@@
The methods used are summarized in Table I. In order to avoid information loss, we add constraints on some transformations, such as limiting the scope of scale transformations, preventing over-dark images. By transforming blocks from different spectral bands within the same spatial region, we increase the diversity of transformations is increased. This forces the model to focus on a broader range of spectral information for correct classification, thereby enhancing the impact of perturbations in the spectral dimension and improving the effectiveness of adversarial examples.
@@
Figure 3: Feature maps of different channels of the image within the same layer of the same model, along with their variances.
III-D Feature Distancing Loss
Current research on adversarial examples for RSI classification rarely considers the image features extracted by the model. However, these features are transferable between models with different architectures [29]. Relying exclusively on information from the classification-layer is inadequate for effectively disrupting the deep-level features of the examples. Therefore, we believe that leveraging more feature information during the attack phase could improve adversarial transferability.
Based on this, we introduce the distance between the feature mappings of the original and adversarial examples as the primary guiding loss. This way, the gradient of the intermediate loss with respect to $x$ can maximize the loss in fewer steps, thereby mitigating the vanishing gradient problem. The goal at this stage is to increase the distance between the features of the original and adversarial examples, thereby minimizing the probability that the feature distribution of the adversarial example belongs to a specific class. As the intermediate feature mapping distance between $x$ and $x_{adv}$ increases, it also leads to a larger distance in the final output of the neural network, thereby increasing the loss function and causing mispredictions [21]. We refer to the process of maximizing feature loss as the distancing operation between image features.
III-D1 Weighted Feature Distancing
@@
When using the feature distancing loss, we can choose to make the features of all layers distant, just like TAP. However, due to differences in model architectures, the early layers of some models fail to learn truly salient features and semantic concepts, while the later layers exhibit strong model-specific features[29]. In contrast, the intermediate layers tend to possess well-separated class representations and are less correlated with the model’s architecture. Therefore, we compute the feature distance of the intermediate layers as the loss. Even when extracting features from the same convolutional layer of a model, significant differences can still be observed between the feature maps of different channels, and the variance of features in each channel also varies (variance is a statistic reflecting the richness of texture information in an image or feature), as shown in Fig. 3. When the variance in a specific channel is large, the paper [36] suggests that the features extracted by that channel are non-robust features related to the true class, which have a greater impact on the model’s classification results.
If the distance is calculated by simply averaging the features across all channels, this variability is overlooked. Therefore, to make the perturbation more effectively disrupt the features within the adversarial example, we calculate the weight of each channel based on the magnitude of its variance, thereby giving non-robust features a greater influence in the loss. This may increase the probability of the adversarial examples deviating from the ground-truth label. Let the intermediate layer $l$ of the substitute model $f$ be denoted as $f^{l}$. Assume $A^{l}\in R^{C_{l}*D_{l}}$represents the features of the adversarial example $x_{adv}$ at layer $l$, where $C_{l}$ is the number of channels and $D_{l}$ is the product of width and height. Similarly, $X^{l}\in{R^{C_{l}*D_{l}}}$ represents the features of the original clean example $x$ at layer $l$. Incorporating the feature weights defined by the variance ratio into the distance function. Let $\sigma_{k}^{X^{l}}$ denote the variance of the $k$-th channel in $X^{l}$, which can be expressed by the following formula:
@@
$\mu_{k}^{X^{l}}=\frac{1}{D_{l}}\sum\nolimits_{i=1}^{{D_{l}}}{X_{k,i}^{l}}$ , $X_{k,i}^{l}$ denotes the $i$-th feature value in the $k$-th channel of $X_{l}$.
Then, we can calculate the weight for each channel, which is the proportion of the variance of a single channel to the total variance across all channels. This can be expressed as follows:
@@
When $W_{k}$ is higher, it shows that the model pays more attention to the $k$-th channel.
For each convolutional layer in a CNN, convolutional kernels extract various features from the input. There is an overall correlation among features across different channels, such as feature distribution, but Euclidean distance ignores such correlations. Furthermore, the Chi-squared distance provides a fair starting point for evaluating the significance of differences between each pair of channels. Therefore, we introduce weights into the Chi-squared distance to compute the feature difference between $x$ and $x_{adv}$. It can be expressed as:
$L_{F}\left({x_{adv},x}\right)$ $=W\cdot d\left(f^{l}\left(x_{adv}\right),f^{l}\left(x\right)\right)$
$=W\cdot d({A^{l}},{X^{l}})$ (4)
@@
We define the above loss as the objective for generating adversarial examples. However, due to the pooling layers in CNNs, some information is inevitably lost, and certain intermediate layer features have lower correlation with the true class of the image [58]. This limitation reduces the attack success rate when relying solely on intermediate layer information. To address this, we integrate output layer information into an auxiliary loss:
@@
Thus, by combining the previous $L_{F}$ with the auxiliary loss, we obtain the total loss, which can be expressed as follows:
@@
where $\eta$ is a hyperparameter representing the proportion of the auxiliary loss.
By using this loss to iteratively update the perturbation, this method makes more comprehensively utilizes the feature information extracted by the model, causing the feature maps of the adversarial examples to move away from the high confidence regions of the true class. Additionally, since different channel features have varying impacts on the classifier’s prediction results, it achieves more effective adversarial perturbations.
III-D2 Enlargement of the original examples features
If we directly increase the feature distance between the original and adversarial examples, the feature values of the adversarial examples at corresponding positions may either increase or decrease. This imbalance in variation could lead to oscillation issues during gradient updates, preventing effective disruption of features related to the true class. Therefore, we employ a feature enlargement method, scaling up the feature values of the original examples by a certain factor to make them larger than those of the initial adversarial examples. In this way, when optimizing based on the target function to achieve the effect of feature distancing, the feature values at the corresponding positions of the adversarial examples should be updated in the direction of decrease. Although transformation operations are applied to adversarial examples, Ding et al. [59] pointed out that data augmentation operations only alter the spatial positions and areas of different feature components in the feature map without changing their numerical relationships. Thus, overlapping regions can still suppress the feature values within adversarial examples. By averaging the gradients of multiple adversarial examples, the perturbation more stably suppresses the features related to the ground-truth label in these adversarial examples, thereby further enhancing the effectiveness of the feature distancing strategy. The updated formula is as follows:
${L_{F}}\left(x_{adv},x\right)$ $=W\cdot d\left({f^{l}}\left(x_{adv}\right),\lambda*{f^{l}}\left(x\right)\right)$ (7)
$=W\cdot d\left({A^{l}},\lambda*{X^{l}}\right)$
where $\lambda$ is the scaling factor for enlarging the feature values of the original examples. Thus, the optimization formula:
$x_{adv}=\arg\max\left({L_{F}}+\eta*{L_{A}}\right),s.t.\left\|\delta\right\|\leq\varepsilon$ (8)
This method not only preserves the diversity introduced by data augmentation but also enhances the destructive capability of adversarial examples.

### v2

@@ III-A Preliminary
In this study, we refer to the known model used to generate adversarial examples as the substitute model, denoted as $f$. The input HSIs are denoted as $x\in X\subset{R^{C\times H\times W}}$, where $H$ and $W$ represent the height and width, respectively, and $C$ is the number of spectral bands. The ground-truth label of $x$ is $y$, and the adversarial example is $x_{adv}\in X\subset R^{C\times H\times W}$. The goal of this study is to find a small perturbation $\delta\subset R^{C\times H\times W}$ that satisfies the constraint condition and enables the adversarial example $x_{adv}=x+\delta$ to maximally mislead the model’s predictions. The loss function is denoted as $L\left(\cdot\right)$. So the process of generating adversarial examples is naturally achieved by maximizing the loss function, which can be expressed as:
${x_{adv}}=\arg\max L\left(f\left(x_{adv}\right),y\right),s.t.{\left\|\delta\right\|_{p}}\leq\varepsilon$ (1)
Typically, the $l_{p}$-norm is used to constrain the perturbation $\delta$ (in this paper, the $l_{1}$-norm is utilized), where $\varepsilon$ is the perturbation budget.
Figure 1: The overall workflow of the proposed method. Before feeding the image into the substitute model, the 3D structure-invariant transformation method is applied to adversarial examples, which divides the examples into blocks in the spatial and spectral dimensions and then uses random transformation operation to each block. While preserving the image structure, it not only helps generate more diverse images but also maintain the global semantic information. Subsequently, the weighted feature divergence loss and cross-entropy classification loss are computed to obtain the total loss, so as to obtain the average gradient, which is iteratively updated to refine the perturbation. Here, $X^{l}$ and $A^{l}$ denote the feature maps of the original and the adversarial examples at layer $l$, respectively.
@@ III-B Overall Process
The pipeline of our method is illustrated in Fig. 1. The transferability of adversarial examples is enhanced through two key steps: 3D structure-invariant transformation and weighted feature divergence loss. Before inputting the adversarial examples into the substitute model, $n$ copies will be generated. These $n$ copies are randomly divided along both the spatial and spectral dimensions. Subsequently, each block in the copy will be randomly applied with a data transformation operation to generate a set of enhanced transformed images, which will serve as input for the substitute model. Then, we calculate the weighted feature divergence loss $L_{F}$ between the original example and the adversarial example, which is combined with the cross-entropy classification loss to form the total loss function. Because the transformed images may differ significantly from the original ones, the calculation of gradients could introduce instability. To solve this problem, we calculate the average gradient according to the average loss of multiple transformed images. Algorithm 1 outlines the perturbation generation process integrated with the baseline method MI-FGSM [20], where $Clip_{x,\varepsilon}$ means to crop the adversarial examples in $x-\varepsilon,x+\varepsilon$]. The details of the two key steps are introduced below.
III-C 3D Structure-Invariant Transformation
Data transformation methods can mitigate the issue of adversarial examples overfitting to the substitute model. However, previous methods usually do not take the image structure into account when transforming the images. This may result in the enhanced image losing important semantic information. Therefore, this paper proposes the structure-invariant method to transform the HSI images when generating adversarial examples. Specifically, given an image $x$, we randomly divide it into blocks and then apply different transformation strategies to each block. In this process, the relative structural relationship between the image blocks can be completely preserved. Since transforming each block does not impair the task such as target recognition [58], this structure-invariant transformation strategy not only helps generate more diverse images but also maintain the global semantic information which are important for downstream tasks.
@@
Figure 2: (a) Schematic of image division. If the spatial dimension division parameter is set to 2, and the spectral dimension is set to 3, then a 2×2×3 block structure will be obtained; (b) Schematic of image transformation. Taking spatial transformation as an example, the image is divided into 2×2 blocks, with different transformations applied to each block as shown.
Since HSIs have rich spatial and spectral information, dividing the HSI can be carried out in both spatial and spectral dimensions. When the structure-invariant transformation is performed only in the spatial dimension, i.e., 2D structure-invariant transformation, the transformation effect will be applied to all spectral channels of each block. Nevertheless, different DNN models usually focus on distinct discriminative spectral bands [32]. Therefore, in order to further improve the diversity of HSIs, we adopt the 3D structure-invariant transformation to HSIs. That is, we randomly divide each image into $a\times a$ blocks in the spatial dimension and $b$ blocks in the spectral dimension, and finally get the three-dimensional block structure of $a\times a\times b$, as shown in Fig. 2. Then, we randomly select different transformation operations from the candidates and apply them to different blocks, thereby generating a set of diverse image variants. The transformation diagram is shown in Fig. 2. The candidate transformation operations used to blocks are summarized in Table I. In order to avoid information loss, we add constraints on some operations, such as limiting the scope of scale transformations, preventing over-dark images.
Fig. 3 shows the comparison between the attention heatmaps of the original example and the transformed examples using the proposed 3D structure-invariant strategy. It can be observed that the model has different attention regions for the images before and after transformation. The use of multiple transformed examples enables the model to focus on a broader range of spectral information during the generation of adversarial examples.
Figure 3: This figure compares the attention maps of the original example and its transformed examples. The first and second rows represent the spatial and spectral dimensions, respectively. The first four columns are the ResNet-18 model, corresponding to the original example and three transformation examples randomly obtained through our proposed transformation method; the last four columns are the same but from the VGG-11 model.
@@
  Input: HSI example $x$, substitute model $f$, loss function $L$, transformation $T\left(\right)$, perturbation $\delta$, perturbation budget $\varepsilon$, step size $\alpha$, decay factor $\mu$, momentum term $m_{i}$, gradients $g_{i}$, number of copies NN, iteration steps II.
@@
3:   Construct a set $X^{adv}_{T}$ of $N$ transformed images using $T()$:
4:      $X^{adv}_{T}=T(x+\delta_{i})*N$
@@
6:   Calculate weighted feature divergence loss according to (4);
7:   Calculate the cross-entropy classification loss according to (5);
@@
10:   Calculate the average gradient of NN copies:
11:     $g_{i}^{ave}=\frac{1}{N}\left(\sum_{j=0}^{N}(\nabla_{\delta}L(x_{j}^{adv},x))\right),(x_{j}^{adv}\in X^{adv}_{T})$
@@
16:   Update adversarial examples: $x_{i+1}^{adv}=Clip_{x,\varepsilon}\{x+\delta_{i}\}$
@@
III-D Weighted Feature Divergence Loss
Current research on adversarial examples for HSI classification rarely considers the image feature maps extracted by the model. However, these features are transferable between models with different architectures [28]. Relying exclusively on information from the classification-layer is inadequate for effectively disrupting the deep-level features of the examples.
Based on this, we introduce the distance between the intermediate feature maps of the original and adversarial examples as the primary guiding loss, and refer to it as the weighted feature divergence loss. By increasing the difference between the feature maps of the original and adversarial examples, our purpose is to effectively perturb the features within adversarial examples so that they deviate from the original feature representations. As there are several layers in the model to extract different level features, we take the feature map of a certain layer for feature loss calculation.
To guide the perturbations to suppress features associated with the true class, it first enlarges the feature values of the original example by a fixed factor, ensuring that the adversarial example’s feature values are lower. Next, it computes the variance of each channel within the feature map to assign different weights to different channels, thereby destroying the features that have a greater impact on HSI classification. We detail the key points of the weighted feature divergence loss calculation as follows.
@@ III-D Weighted Feature Divergence Loss
Figure 4: Feature maps of different channels in the same layer of the VGG-11 model, along with their variances.
III-D1 Enlarge the Feature Values of Original Examples
When calculating the feature distance, if we directly increase the difference between the feature maps of the original and adversarial examples, the perturbation during the update process may cause the values in the high-activation regions of the adversarial example’s features to either increase or decrease, which will not effectively suppress the features related to the true class.
To address this issue, we enlarge the original features by a certain factor as shown in Eq.(4), so that the values in the high-activation regions of the initial adversarial features are significantly lower than those in the original example. In this way, when updating the perturbation based on the weighted feature divergence loss, the values in the high-activation regions of the adversarial example’s features related to the true class will be more likely to decrease.
Although transformation operations are applied to adversarial examples, Ding et al. [59] pointed out that data augmentation operations only alter the spatial positions and areas of different feature components in the feature maps. Thus, neighboring or overlapping regions can still suppress the feature values within adversarial examples. By averaging the gradients of multiple adversarial examples, the perturbation more stably suppresses the features related to the ground-truth label in these adversarial examples, thereby enhancing the effectiveness of the weighted feature divergence strategy.
III-D2 Weighted Feature Divergence
@@
Even features extracted from the same convolutional layer of a model, differences can still be observed between the feature maps of different channels, and the variance of features in each channel also varies, as shown in Fig. 4. When the variance in a specific channel is large, the features extracted by that channel are non-robust features related to the true class [33], which have a greater impact on the model’s classification results.
If we simply averages these channels when computing the distance between features, the differences between the channels are ignored. Therefore, to make the perturbation more effectively disrupt the non-robust features within the adversarial example, we calculated the weights of each channel according to their variance in the feature maps, and give non-robust features a greater influence in the loss. This increases the probability of the adversarial examples deviating from the ground-truth label.
Let the intermediate layer $l$ of the substitute model $f$ be denoted as $f^{l}$. $X^{l}\in{R^{C_{l}*D_{l}}}$ represents the feature maps of the original example $x$ at layer $l$, where $C_{l}$ is the number of channels of the feature maps and $D_{l}$ is the product of width and height of the feature maps. Let $\sigma_{k}^{X^{l}}$ denote the variance of the $k$-th channel in $X^{l}$, which can be expressed by the following formula:
@@
[…]
Then, the weight of each channel is defined as the proportion of the variance of a single channel to the total variance across all channels, which can be expressed as follows:
@@
When $W_{k}$ is higher, the model pays more attention to the $k$-th channel.
For each convolutional layer in DNN, convolutional kernels extract various feature maps from the input. There is an overall correlation among features across different channels, such as feature distribution, but Euclidean distance ignores such correlations. In contrast, the Chi-squared distance provides a fair starting point for evaluating the differences between each pair of channels. Let $A^{l}\in R^{C_{l}*D_{l}}$ represents the features of the adversarial example $x_{adv}$ at layer $l$, where $C_{l}$ is the number of channels and $D_{l}$ is the product of width and height. With the Chi-squared distance, the weighted feature divergence loss, i.e., the weighted feature maps difference between $x$ and $x_{adv}$, can be expressed as:
$L_{F}\left({x_{adv},x}\right)$ $=W\cdot d\left(f^{l}\left(x_{adv}\right),\lambda f^{l}\left(x\right)\right)$
$=W\cdot d({A^{l}},\lambda{X^{l}})$ (4)
@@
where $\lambda$ is the scaling factor for enlarging the feature values of the original examples.
Moreover, due to the pooling layers in DNN models, some information is inevitably lost, and certain intermediate layer feature maps have lower correlation with the true class of the image [60]. This limitation reduces the attack success rate when relying solely on intermediate layer information. To address this, the cross-entropy classification loss which is derived according to the information of the output layer is also adopted, which is defined as:
@@
where $CE(\cdot)$ represents the cross-entropy function. Thus, by combining the previous $L_{F}$ with the cross-entropy classification loss, the total loss of the proposed method is defined as Eq.(6).
@@
where $\eta$ is a hyperparameter representing the proportion of the cross-entropy classification loss.
The final optimization formula can be expressed:
$x_{adv}=\arg\max\left({L_{F}}+\eta*{L_{A}}\right),s.t.\left\|\delta\right\|_{1}\leq\varepsilon$ (7)
We greatly enhanced the diversity of the input samples through the use of 3D structure-invariant transformation techniques. This compels the model’s attention to extend from local features to multiple global regions, thereby establishing richer discriminative zones in the feature space. This provides the weighted feature divergence loss with more optimization directions and richer feature information, while the loss itself also reinforces the utilization of the feature space. Together, these two factors strengthen the adversarial transferability.
Refer to captionRefer to captionRefer to caption
(a) HoustonU 2018
(b) PaviaU
(c) Indian Pines
Figure 5: False color images and ground-truth maps of the (a) HoustonU 2018, (b) PaviaU, and (c) Indian Pines datasets.
TABLE II: Training and testing sets of the HoustonU 2018, PaviaU, and Indian Pines datasets
HoustonU 2018 PaviaU Indian Pines
Class Train Test Class Train Test Class Train Test Class Train Test Class Train Test
1 5000 4799 11 1000 24002 1 900 5731 1 20 26 11 1200 1255
2 10000 22502 12 100 516 2 900 17749 2 700 728 12 300 293
3 500 184 13 10000 36358 3 900 1199 3 400 430 13 100 105
4 10000 3588 14 5000 4849 4 900 164 4 120 117 14 630 635
5 3000 2048 15 5000 1937 5 900 445 5 240 243 15 200 183
6 3000 1516 16 10000 1475 6 900 4129 6 360 370 16 50 43
7 200 66 17 100 49 7 900 430 7 15 13
8 10000 29762 18 3000 3578 8 900 2782 8 230 248
9 10000 213684 19 3000 2365 9 900 47 9 10 10
10 10000 35810 20 3000 3824 10 500 472